Circular No. 5873

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

HD 98800 D. Bockelée-Morvan, P. Colom, J. Crovisier, D. Gautier and E. Gérard,

Observatoire de Paris-Meudon; T. Owen, University of Hawaii; and D. Despois, Observatoire de Bordeaux, report: “A search was made with the Nançay radio telescope (beam 3′.5 × 19′) for the 1667-MHz transition of the OH radical toward the dusty K5V star HD 98800 (Zuckerman and Becklin 1993, Ap.J. 406, L33; see also IAUC 5804). An absorption signal detected at a 7σ level on the line area in July may be associated with the star. The peak flux density is –6 mJy, and the central LSR velocity and width of the line are –20 km/s and 30 km/s, respectively. The line is displaced toward the blue by 26 km/s with respect to the velocity of the star. No signal was detected at 3′.5 east and west of the nominal position. Both the line characteristics and the galactic coordinates suggest that an interstellar origin is unlikely. Observations at 18 cm with broad-band filters were not able to detect point-source continuum emission above 50 mJy. Further observations are needed to establish firmly whether the detected line is associated with the circumstellar disk around HD 98800.”

NOTICE TO CONTRIBUTORS AND SUBSCRIBERS Effective with IAUC 5874, there will be an increase in the line-charge

rate for those items published on these Circulars that are subject to these charges. The new charge will be $17.50 per line (in the printed version) plus $45.00 per item (the latter reduced to $35.00 for shared items). The items subject to line charges are generally those relating to other than discovery and initial confirmation of comets, unusual minor planets (most of which will now be covered instead on the new Minor Planet Electronic Circulars anyway), novae and supernovae. Continuing astrometric and orbital data on comets (and minor planets), and initial spectroscopic data on novae and supernovae, are also exempt, as are contributions (not all of which are accepted for publication) by amateur astronomers.

Following the responses to the solicitation on IAUC 5735, the Central Bureau has decided to discontinue its telegram (cablegram, telex and mailgram) service (although it can still receive material by telegram). Any credit remaining in the accounts of telegram subscribers will be tranferred to the corresponding Circulars accounts at the end of October.

Effective immediately, those who wish to purchase the Catalogue of Cometary Orbits (cf. IAUC 5792) in printed or diskette form should add $4.00 for postage and handling (or add $10.00 for airmail delivery).

1993 September 30 Brian G. Marsden